OpenStudy (anonymous):

How much work is required to lift a 10N weight from 4.0 m to 40m above the surface of the earth

OpenStudy (cybershadow):

see we got the force F = 10 N we need to move it to the distance dx = 40 -4.0 = 36 m also work = F dx = 10 * 36 = 360 J

OpenStudy (anonymous):

to lift the weight work shud be done against gravity and the force mg needs to be overcome . displacement=40-4=36m work done =forcex displacement=weight x 36=10x36=360N

OpenStudy (anonymous):

units for work are N - m = J